# Nonlinear Impulse Pattern Formulation dynamical social and political prediction algorithm for city planning and public participation

### v1

@@ Abstract
A nonlinear-dynamical algorithm for city planning is proposed as an Impulse Pattern Formulation (IPF) for predicting relevant parameters like health, artistic freedom, or financial developments of different social or political stakeholders over the cause of a planning process. The IPF has already shown high predictive precision at low computational cost in musical instrument simulations, brain dynamics, and human-human interactions. The social and political IPF consists of three basic equations of system state developments, self-adaptation of stakeholders, two adaptive interactions, and external impact terms suitable for respective planning situations. Typical scenarios of stakeholder interactions and developments are modeled by adjusting a set of system parameters. These include stakeholder reaction to external input, enhanced system stability through self-adaptation, stakeholder convergence due to mediative interaction adaptation, as well as complex dynamics in terms of direct stakeholder impacts. A workflow for implementing the algorithm in real city planning scenarios is outlined. This workflow includes machine learning of a suitable set of parameters suggesting best-practice planning to aim at the desired development of the planning process and its output.
@@ I Introduction
"We are already colourful" A statement that does not seem to fit into the scientific context and yet has led to the introduction of an algorithm in urban planning to test, question or confirm this statement, for a review see Abas2023 . The statement of the citizens who see themselves exposed to the plans for a future change in their neighbourhood. How can urban planners advance their projects in a targeted manner in the interests of the community and expect a "Yes please, let’s examine the parameters for a liveable future together" instead of a "No thanks, we’ve already set up our lives"? Many approaches have been proposed mitigating this problem, often involve machine learning algorithms.
In times of scarcity of space and land, diverging political goals such as the net zero land strategy and alliances in favour of building even more housing, the smallest parameters that are not taken into account can have a decisive influence on the already fragile balance and even prevent planning altogetherCharturvedi2021 . Active citizens’ initiatives can lead to referendums and long-term uncertainties in urban planning. On both sides of the process, resources and lifetimes are spent fighting against their own system. There is no question that planning in an increasingly dense city is becoming more complex and planning for the future more uncertain. But building takes time and what is built is there for the time being: it creates space and thus shapes society. Both the new neighbourhood and the old neighbourhood need to be well planned or replanned, taking into account all known parameters, so that stable, good communities and thus stable societies can form. With forward-looking planning that takes environmental concerns, demographic aspects and crises into account, an equal society is conceivable.Wilkinson2024 .
@@
Social polarization has often been studied using different nonlinear dynamical methods. Using a Hamiltonian approach of connected social network coherence minimization, a phase transition was found when the amount of connectivity in the society rises above a certain level, which only has happened through extensive internet connectivity over the last decadesMinhPham2020 . This model allows only a polar decision regarding friend/enemy.
@@
The Impulse Pattern Formulation (IPF) model proposed in the present paper to arrive at this aim has first been developed for musical instrument applicationsBader2013 ; Linke2021b ; Linke2019b . It has also been proven to predict human-human interaction in rhythm perception and productionLinke2021a . The IPF has also proven to work in brain dynamicsBader2022 , explaining fundamental aspects like the excitatory/inhibitory neuron concentration or typical brain time scales. The brain IPF also corresponds to EEG measurements of brain synchronization as the physiological cause of musical large-scale tension dynamicsBader2024 .
The IPF takes a viewpoint neuron of, e.g., neurons or musical instrument parts from which an impulse is sent out to several other neurons, musical instrument parts, or any object. This impulse is processed, damped, and returns to the viewpoint objectBader2013 ; Linke2019b ; Bader2021 ; Linke2021b leading to convergent, bifurcating, chaotic, or diverging behavior. Although often modeling a system with only very few nodal points, the IPF has already been shown to be highly precise.
As the IPF already assumes a set of standpoints, which in the case of city planning are stakeholders, and can simulate very different types of temporal stakeholder developments as well as sudden phase changes, converging, bifurcating, and chaotic behavior, it is used as a basis for a political and social prediction process.
In the frappant case discussed above, six such stakeholders were identified. These are the artists, local stores, local residents, residents of Altona, the investor IKEA, and the City of Hamburg. Although the present paper outlines the basic framework of the model discussing real-world cases in detail is beyond its scope. Still, the number of stakeholders used in the model is fixed to six throughout.
Such an iterative, nonlinear-dynamical process is scale-free, capable of modeling sudden phase changes, and includes convergent, bifurcating, complex, and chaotic states. Scaling is a problem in social sciences in general, leading to so-called voltage drops when evidence found in one social group is scaled up to a broader levelList2024 .
@@
The interaction between stakeholders in the model is formulated in two different versions: a direct and a mediated interaction. Therefore, the structural balance theory discussed above is not used as a model input but might emerge from it.
The IPF is an algorithm emerging from a Physical Culture TheoryBader2021 , assuming no difference between humans, nature, culture, or technology by finding all to be electromagnetic fields when broken down to a physical basis. This includes consciousness, which is found to be a tempo-spatial electromagnetic field complex enough to arrive at conscious content of all kinds. Returning to this introductions beginning, where the self-organization nature of life in general was discussed, in this view, cultural artifacts like musical instruments, which are self-organizing structures by their very nature but made by humans, are viewed as enlarging the human body and mind with interacting artifacts as a way to maintain life by reducing a system entropy.
@@ II Method
The IPF furthermore assumes a self-reference of each standpoint upon itself using a self-adaptation variable $\alpha^{s}_{t}$, adapting according to the standpoint itself. The damping variable $\alpha^{s}$ determines the system is stable, bifurcating, chaotic, and unconverting.
@@ II.1 Social IPF system state equations
The general form of the social IPF is
$g^{s}_{t+1}=g^{s}_{t}-\log_{e}\left(\frac{1}{\alpha^{s}}\ g^{s}_{t}-\sum_{i=1}^{N,i\neq s}\beta^{s,i}e^{g^{s}_{t}-e^{s}_{t-d^{s,i}}}\right)\ +I^{s}_{t}.$ (1)
The time delay parameter $d^{s,i}$ it takes until the state $g^{i}$ is acting on $\beta^{s,i}$ might vary according to the relation between stakeholders and therefore need to be a variable.
@@ II.3 Stakeholder interaction equation
The impact of stakeholder $i$ on stakeholder $s$ is according to the interaction variable $\beta^{s,i}$. Changing $\beta^{s,i}$ during modeling is well known from the literature (e.g., Linke2019a ; Linke2021a ). As shown in Fig. 1, a sudden change of $\beta^{s,i}$ directly affects the system state $g$. The system may (after a short transient) converge to a new stable limit. However, it is also possible that the state will change to bifurcations, chaos, and sometimes even diverging states.
Figure 1: different time series of $g$ when suddenly changing beta for an IPF with $N=2$ stakeholders / refelction points: $\alpha=0.369$ is constant. $\beta$ always starts at $0.2$ and is suddenly changed to the different displayed values after 50 iteration steps
Nevertheless, as shown in Fig. 2, continuously changing $\beta^{s,i}$ keeps the system close to its fixed point. Thus, the system appears to be stable even in regions where the fixpoint is not. However, as soon as $\beta^{s,i}$ does not change anymore, the system gradually shows the expected behavior.
Figure 2: different time series of $g$ when continously changing beta for an IPF with $N=2$ stakeholders / refelction points: $\alpha=0.369$ is constant. $\beta$ always starts at $0.2$ and is linearly changed to the different displayed values between the 40th and the 60th iteration step
The mutually changing of $\beta^{s,i}$ of different stakeholders can be subject to several interaction processes. In direct interaction, the interaction variable is proportional to the system variable of the impacting stakeholder $i$ multiplied by a constant $W^{x,i}$ like
@@
We refer to this version as direct impact in the following.
@@
$\beta^{s,i}_{t+1}=\beta^{s,i}_{t}-V^{s,i}\left(g^{s}_{t}-g^{i}_{t-d^{s,i}}\right)\ .$ (6)
This accounts for the fact that stakeholders are in constant interaction and try to balance their impact constantly in time from one to another. We refer to this version as mediating impact in the following.
So $\beta^{s,i}_{t}$ is the third varying term to be decided for in a social scenario.
The time delay can be formulated more generally like
@@
Here, the system state $g^{s}$ is not compared with its present state to a previous state of $g^{i}$ at $d^{s,i}$ but also in its previous state $d^{s,i}$. This would imply that one stakeholder remembers a previous state or that the impacting stakeholder i refers to a previous stake of s. This is a more general version to be included, although it is not expected to be the general case. Still, we will discuss the impact of such a case below in some detail, too.
@@ II.4 Workflow of the social, political, and city planning IPF
Tab. 1 gives an overview of the varying terms, again as discussed above. There might or might not be an external impact (Varying Term I) for each stakeholder $s$ separately. There might or might not be self-adaptation (Varying Term II) for each stakeholder separately. And interaction might be direct or mediated (Varying Term III). This last condition might not be exclusive, and both interactions might take place at the same time. In this paper, we first consider both terms alone to discuss their properties and leave a combined version for future studies.
The fixed parameters of the model are summarized in Tabl. 2. The self-adaptation weights determine the self-adaptation strength of the variable $\alpha^{s}_{t}$, as are the direct and mediating interaction weight for the interaction variables $\beta^{s,i}_{t}$. The delays of stakeholders interacting with each other are fixed parameters.
@@
Variation I External impact Self-adaptation Direct interaction
var. $\alpha^{s}_{t+1}=\left(1/\alpha^{s}_{t}-C^{s}(g^{s}_{t}-g^{s}_{t-1})\right)^{-1}$ $\beta^{s,i}_{t+1}=W^{s,i}g^{i}_{t-d^{s,i}}$
Variation II No external impact No self-adaptation Mediating interaction
@@
Self-adaption weight $C^{s}$
Direct interaction weight $W^{s,i}$
Mediating interaction weight $V^{s,i}$
@@ II.4.7 Interpreting social or political actions from predicted parameters
The parameters of self-adaptation weight $C^{s}$ and direct or mediated interaction weight $W^{s,i}$ or $V^{s,i}$ respectively, need to be translated into action or decisions.
With best-fit results in the case of mediated interactions between several stakeholders demanding a high weight, stakeholders need to make sure both sides interact closely to achieve the appropriate goal. This might also determine a best-fit interaction delay, which could imply regular meetings, financial benefits, a suggested time interval for artistic or cultural activity, or the like.

### v2

# Nonlinear dynamical social and political prediction algorithm for city planning and public participation using the Impulse Pattern Formulation

@@ Abstract
A nonlinear-dynamical algorithm for city planning is proposed as an Impulse Pattern Formulation (IPF) for predicting relevant parameters like health, artistic freedom, or financial developments of different social or political stakeholders over the cause of a planning process. The IPF has already shown high predictive precision at low computational cost in musical instrument simulations, brain dynamics, and human-human interactions. The social and political IPF consists of three basic equations of system state developments, self-adaptation of stakeholders, two adaptive interactions, and external impact terms suitable for respective planning situations. Typical scenarios of stakeholder interactions and developments are modeled by adjusting a set of system parameters. These include stakeholder reaction to external input, enhanced system stability through self-adaptation, stakeholder convergence due to adaptive interaction, as well as complex dynamics in terms of fixed stakeholder impacts. A workflow for implementing the algorithm in real city planning scenarios is outlined. This workflow includes machine learning of a suitable set of parameters suggesting best-practice planning to aim at the desired development of the planning process and its output.
@@ I Introduction
"We are already colourful" A statement that does not seem to fit into the scientific context and yet has led to the introduction of an algorithm in urban planning to test, question or confirm this statement, for a review see Abas2023 . The statement of the citizens who see themselves exposed to the plans for a future change in their neighbourhood. How can urban planners advance their projects in a targeted manner in the interests of the community and expect a "Yes please, let’s examine the parameters for a liveable future together" instead of a "No thanks, we’ve already set up our lives"? Many approaches have been proposed mitigating this problem, often involving machine learning algorithms.
In times of scarcity of space and land, diverging political goals such as the net zero land strategy and alliances in favour of building even more housing, the smallest parameters that are not taken into account can have a decisive influence on the already fragile balance and even prevent planning altogetherCharturvedi2021 . Active citizens’ initiatives can lead to referendums and long-term uncertainties in urban planning. On both sides of the process, resources and lifetimes are spent fighting against their own system. There is no question that planning in an increasingly dense city is becoming more complex and planning for the future more uncertain. But building takes time and what is built is there for the time being: it creates space and thus shapes society. Both the new neighbourhood and the old neighbourhood need to be well planned or replanned, taking into account all known parameters, so that stable, sustainably communities and thus stable societies can form. With forward-looking planning that takes environmental concerns, demographic aspects and crises into account, an equal society is conceivable.Wilkinson2024 .
@@
Social polarization has often been studied using different nonlinear dynamical methods. Using a Hamiltonian approach of connected social network coherence minimization, a phase transition was found when the amount of connectivity in the society rises above a certain level, which only has happened through extensive internet connectivity over the last decadesMinhPham2020 . This model allows only a polar decision regarding friend/enemy. The role of tipping points has been stressed in a model of hysteresis loop polaization processesMacya2022 . This model is agent-based rather than using differential equations. Another approach is using statistical methods like Boltzman-type dynamicsLoy2022 . This allows for analytical descriptions of phase-changes in polarization processes.
@@
The Impulse Pattern Formulation (IPF) model proposed in the present paper to arrive at this aim has first been developed for musical instrument applicationsBader2013 ; Linke2021b ; Linke2019b . It has also been proven to predict human-human interaction in rhythm perception and productionLinke2021a , according to previous physiologicalBuzsaki2006 Ha2017 and model findingsFuhrmann2002 Fardet2020 . The IPF has also proven to work in brain dynamicsBader2022 , explaining fundamental aspects like the excitatory/inhibitory neuron concentration or typical brain time scales. The brain IPF also corresponds to EEG measurements of brain synchronization as the physiological cause of musical large-scale tension dynamicsBader2024 .
The IPF takes a viewpoint neuron of an object from which an impulse is sent out to several other neurons, musical instrument parts, or any object. This impulse is processed, damped, and returns to the viewpoint objectBader2013 ; Linke2019b ; Bader2021 ; Linke2021b leading to convergent, bifurcating, chaotic, or diverging behavior. Although often modeling a system with only very few nodal points, the IPF has already been shown to be highly precise.
As the IPF already assumes a set of standpoints, which in the case of city planning are stakeholders, and can simulate very different types of temporal stakeholder developments as well as sudden phase changes, converging, oscillating, bifurcating, and chaotic behavior, it is used as a basis for a political and social prediction process.
In the frappant case discussed at the beginning of the Introduction, six such stakeholders were identified. These are the artists, local stores, local residents, residents of Altona, the investor IKEA, and the City of Hamburg. Although the present paper outlines the basic framework of the model discussing real-world cases in detail is beyond its scope. Still, the number of stakeholders used in the model is fixed to six throughout.
Such an iterative, nonlinear-dynamical process is scale-free, capable of modeling sudden phase changes, and includes convergent, oscillating, bifurcating, complex, and chaotic states. Scaling is a problem in social sciences in general, leading to so-called voltage drops when evidence found in one social group is scaled up to a broader levelList2024 .
@@
The interaction between stakeholders in the model is formulated in two different versions: a fixed and an adaptive interaction. Therefore, the structural balance theory discussed above is not used as a model input but might emerge from it.
The IPF is an algorithm emerging from a Physical Culture TheoryBader2021 , assuming a common ground for humans, nature, culture, or technology by finding all to be electromagnetic fields when broken down to a physical basis. This includes consciousness, which is found to be a tempo-spatial electromagnetic field complex enough to arrive at conscious content of all kinds. Returning to this introductions beginning, where the self-organization nature of life in general was discussed, in this view, cultural artifacts like musical instruments, which are self-organizing structures by their very nature but made by humans, are viewed as enlarging the human body and mind with interacting artifacts as a way to maintain life by reducing a system entropy.
@@ II Method
The IPF furthermore assumes a self-reference of each standpoint upon itself using a self-adaptation variable $\alpha^{s}_{t}$, adapting according to the standpoint itself. The damping variable $\alpha^{s}$ determines whether the system is stationary, oscillating, bifurcating, chaotic, and nonconverging.
In the follow, stationary time series are those with a stationary fixed point the system has converged to. Limit cycles lead to oscillatory behaviour. Bifurcations are present in cases of alternating values of the time series of different order. The lowest order bifurcations are those a time series alternates between two different, adjacent values, higher-order bifurcations are present when alternative values are more then two, typically four or eight. Finally, chaotic time series are highly complex, still not perfectly random.
This terminology has shown useful in musical applications, where a stationary time series corresponds to a musical sound of constant pitch, limit cycles are alternating pitches, so musical vibrato, while bifurcations are so-called multiphonics, where two are more musical pitches are present at the same time. Finally, chaotic behaviour is present during initial transients as well as in certain types of musical articulation like with low-pressure wind or bowed instrument sound production.
@@ II.1 Social IPF system state equations
The Impulse Pattern Formulation (IPF) was first motivated to describe sound propagation in musical instrumentsBader2013 Linke2019a . There, impulse-like shaped acoustic waves are emerging from strings or reeds which travel through the instrument, are reflected and damped, and return to their point of origin. In reed instruments a pressure impulse is creased at the mouthpiece, travels through the reed, is reflected at the bell, is damped through radiation and wall losses, and returns to the reed again to trigger a new impulse. With stringed instruments, the force of the string upon a soundboard is impulse-shaped too. It travels through plates, ribs or air volumes, is damped along the way only to return to the string point. It can be shown that only through this nonlinear dynamical behaviour of such systems a constant periodicity of the system appears which is the musical pitch.
Such a system can be described in its most simple form like
$e^{-\frac{\partial g}{\partial t}}=\frac{1}{\alpha}g\ ,$ (1)
with system parameter $g$, the impulse strength, and damping $\alpha$. The exponential models the damping of such a system, which is typically exponentially decaying in energy. In this formulation the precise shape of the impulse is not taken into consideration but only its strength is represented as system parameter g. Also, the IPF is taking a viewpoint, the point the impulse is sent out and returns to.
As impulses leave a viewpoint and return there at discrete time points. Therefore, the system state $g_{t}$ is introduced like
$e^{-\frac{g_{t+1}-g_{t}}{\Delta t}}=\frac{1}{\alpha}g_{t}\ ,$ (2)
for adjacent times points t. Solving for $g_{t+1}$ leads to the iterative equation
$g_{t+1}=g_{t}-\log_{e}{\frac{g_{t}}{\alpha}}\ .$ (3)
Assuming systems with multiple reflection point, this equation can be written in a more general form which directly can be used as the social IPF like
$g^{s}_{t+1}=g^{s}_{t}-\log_{e}\left(\frac{1}{\alpha^{s}}\ g^{s}_{t}-\sum_{i=1}^{N,i\neq s}\beta^{s,i}e^{g^{s}_{t}-g^{s}_{t-d^{s,i}}}\right)\ +I^{s}_{t}.$ (4)
Here $g^{s}_{t}$ is now defined for all stakeholders $s=1,2,3,...N$ in the model. A time delay parameter $d^{s,i}$ with $i=1,2,3,...N$ and $i\neq s$ is necessary to define the time it takes until a stakeholder $i$ is acting on a stakeholder $s$. The stakeholder interaction strength $\beta^{s,i}$ again of stakeholder $i$ acting on stakeholder $s$ might vary according to the relation between stakeholders and therefore need to be a variable.
@@ II.3 Stakeholder interaction equation
The impact of stakeholder $i$ on stakeholder $s$ is according to the interaction variable $\beta^{s,i}$. Changing $\beta^{s,i}$ during modeling is well known from the literature (e.g., Linke2019a ; Linke2021a ). As shown in Fig. 1, a sudden change of $\beta^{s,i}$ directly affects the system state $g$.
When keeping $\beta^{s,i}=0.2$ (green) the stationary state is kept. Increasing $\beta^{s,i}$, after a short oscillatory transient phase $g^{s,i}$ converges to a new stationary limit (pink). Still, when decreasing $\beta^{s,i}$ to 0.1 a low-order bifurcation state is reached where $g1{s,i}$ alternates between two values (red). Decreasing $\beta{s,i}$ to zero leads to a chaotic behaviour (blue).
Figure 1: Time series of $g^{s}$ with $N=2$ interacting stakeholders and constant $\alpha=0.369$ when suddenly changing $\beta^{s,i}$ at iteration step 51. $\beta^{s,i}$ always starts at $0.2$ and is suddenly increased to 0.25 (pink) converging to a new stationary limit or decreased to 0.1 or 0 leading to a low-order bifurcation or a chaotic behaviour respectively.
Nevertheless, as shown in Fig. 2, continuously changing $\beta^{s,i}$ keeps the system close to its fixed point. Thus, the system appears to be stationary in the sense of not displaying bifurcations or chaos even in regions where the fixed point is not stationary. However, as soon as $\beta^{s,i}$ does not change anymore, the system gradually shows the expected behavior.
Figure 2: Time series of $g^{s}$ with $N=2$ interacting stakeholders and constant $\alpha=0.369$ when gradually changing $\beta^{s,i}$ between iteration step 40 to 60. Start and end values of $\beta^{s,i}$ are the same as in Fig. 1 and the same system states are reached in the end. Still, during transition $g^{s,i}$ is stationary even if the fixed point is not.
The mutually changing of $\beta^{s,i}$ of different stakeholders can be subject to several interaction processes. In fixed interaction, the interaction variable is proportional to the system variable of the impacting stakeholder $i$ multiplied by a constant $W^{x,i}$ like
@@
We refer to this version as fixed impact in the followingBerner2023 Berner2021 .
@@
$\beta^{s,i}_{t+1}=\beta^{s,i}_{t}-V^{s,i}\left(g^{s}_{t}-g^{i}_{t-d^{s,i}}\right)\ ,$ (9)
where $V^{s,i}$ is a constant analog to $W^{s,i}$.
This accounts for the fact that stakeholders are in constant interaction and try to balance their impact constantly in time from one to another. We refer to this version as adaptive impact in the following.
So $\beta^{s,i}_{t}$ is the third term to be decided for, therefore called Varying Term II below in a social scenario.
The time delay can be formulated alternatively like
@@
in its present state with a previous state of $g^{i}$ delayed by $d^{s,i}$ but in its previous state delayed by $d^{s,i}$
Here, the system state $g^{s}$ is not compared in its present state with a previous state of $g^{i}$ delayed by $d^{s,i}$ but in its previous state delayed by $d^{s,i}$. This would imply that one stakeholder remembers a previous state or that the impacting stakeholder i refers to a previous state of s. This is a more general version to be included, although it is not expected to be the general case. Still, we will discuss the impact of such a case below in some detail, too.
@@ II.4 Workflow of the social, political, and city planning IPF
Tab. 1 gives an overview of the varying terms, again as discussed above. There might or might not be an external impact (Varying Term I) for each stakeholder $s$ separately. There might or might not be self-adaptation (Varying Term II) for each stakeholder separately. And interaction might be fixed or adaptive (Varying Term III). This last condition might not be exclusive, and both interactions might take place at the same time. In this paper, we first consider both terms alone to discuss their properties and leave a combined version for future studies.
The fixed parameters of the model are summarized in Tabl. 2. The self-adaptation weights determine the self-adaptation strength of the variable $\alpha^{s}_{t}$, as are the fixed and adaptive interaction weights for the interaction variables $\beta^{s,i}_{t}$. The delays of stakeholders interacting with each other are fixed parameters.
@@
Version I External impact Self-adaptation Fixed interaction
variable $\alpha^{s}_{t+1}\left(1/\alpha^{s}_{t}-C^{s}(g^{s}_{t}-g^{s}_{t-1})\right)^{-1}$ $\beta^{s,i}_{t+1}=W^{s,i}g^{i}_{t-d^{s,i}}$
Version II No external impact No self-adaptation Adaptive interaction
@@
Self-adaptation weight $C^{s}$
Fixed interaction weight $W^{s,i}$
Adaptive interaction weight $V^{s,i}$
@@ II.4.1 Decision on stakeholders
Stakeholder< s Interest Data/Measurements $g^{s}_{t}$
Artists Artistic freedom qm rehearsal and performance spaces
Residents at Frappant Health condition Volume of traffic
Residents in neighbourhood Attractive shopping area Number of visitors (but/train users)
Local stores Increasing Rents/Take profits Rent / financial balance
Administration/Government Social/financial balance elections
Investors Take profits financial balance
Table 3: List of stakeholders of Frappant example. Each stakeholder has an interest concerning the city planning project. This interest is transferred into quantitative figures like financial balance for investors or qm of performance or rehearsal space for artists over time as $g^{s}_{t}$ in the IPF.
Tab. 3 lists the stakeholders, their interest in the city planning project and the data or measurements used as quantitative $g^{s}_{t}$ values in the model. So e.g. investors are interested in the financial balance, residents of Frappant are interested in their health condition which is threatened due to increased traffic, or artists fear they loose their performance and rehearsal spaces.
All these figures are present as numbers and found in city data bases over the modeling time of the project and with a certain discretization time step. The highest resolution, e.g. traffic measured on a monthly base, is taken as smallest discretization time step.
Alternatively, the model starts with certain values and calculates these data over time into the future.
@@ II.4.7 Interpreting social or political actions from predicted parameters
The parameters of self-adaptation weight $C^{s}$ and fixed or adaptive interaction weight $W^{s,i}$ or $V^{s,i}$ respectively, need to be translated into action or decisions.
With best-fit results in the case of adaptive interactions between several stakeholders demanding a high weight, stakeholders need to make sure both sides interact closely to achieve the appropriate goal. This might also determine a best-fit interaction delay, which could imply regular meetings, financial benefits, a suggested time interval for artistic or cultural activity, or the like.